OpenStudy (sharkfinsoup):

What is the value of the discriminant in the equation shown below? x^2-6x+4=0

OpenStudy (mathmale):

Compare your equation to ax^2 + bx + c = 0. What are the values of a, b and c? a=? b=? c=?

OpenStudy (sharkfinsoup):

a=x^2 b=-6x c=4

OpenStudy (mathmale):

a, b and c are called "coefficients." Focus on the 1 in 1x^2, the -6 in -6x, and the 4. a=1 b=-6 c=4. Again, do not include the powers of x: x^2, x, x^0. Now, what is the value of b^2?

OpenStudy (sharkfinsoup):

36

OpenStudy (mathmale):

That's right: (-6)^2=36. What is the value of a*c? (a times c)

OpenStudy (sharkfinsoup):

4

OpenStudy (mathmale):

Good! Now please write out "b^2 - 4ac," using your previous calculations. b^2-4ac= ?

OpenStudy (sharkfinsoup):

36-16=20

OpenStudy (mathmale):

You're doing great. That 's the "discriminant" of the given equation. You're done. When told to find the "discriminant," calculate b^2-4ac.

OpenStudy (sharkfinsoup):

oh ok

OpenStudy (sharkfinsoup):

so 20 is the answer

OpenStudy (mathmale):

Were you to continue, to find the two roots of the given equation, you'd evaluate\[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (mathmale):

Yes, the discriminant here is 20, and you're done. But be sure you understand that b^2-4ac is part of the quadratic formula, given above, by which we find the "roots" or "zeros" of the given quadratic equation.

OpenStudy (sharkfinsoup):

Thank you so much for the explanation

OpenStudy (mathmale):

Best to know what you can do with a result such as your discriminant, '20.' You're welcome!